Finding the reaction order from a table...?

So, I'm trying to find the reaction order with respect to the concntration of ClO- and then calculate the full rate law, but I'm a little lost. Here's the table:

Trial [Dye] (M) [ClO-] (M) Initial Rate (M/s)
1 6.00*10^-6 0.152 0.0015618
2 6.00*10^-6 0.303 0.0028939
3 6.00*10^-6 0.455 0.003382
4 6.00*10^-6 0.606 0.0038187

I've gotten different answers depending on the rows I've used. So, using the first and second rows, I calculated that the reaction order is 1. But for the second and third rows, I got about 2. Are they supposed to be different?

Thank you in advance for any help!
Answer #1

The reaction is independent of the concentration of the dye.

Let the rate of the reaction be given as

Rate = k*[ClO-]a

where k = rate constant for the reaction and a = order of the reaction with respect to ClO-.

Trail 1

[ClO-] (M)

Initial Rate (M/s)

Ratio of

Ratio of

1

0.152

0.0015618

-

-

2

0.303

0.0028939

R2/R1 = (0.0028939)/(0.0015618) = 1.8529 ≈ 2.0

(0.303)(0.152) = 1.9934 ≈ 2.0

3

0.455

0.0033820

R3/R1 = (0.0033820)/(0.0015618) = 2.1654 ≈ 2.0

(0.455)/(0.152) = 2.9934 ≈ 3.0

4

0.606

0.0038187

R4/R1 = (0.0038187)/(0.0015618) = 2.4451 ≈ 2.0

(0.606)/(0.152) = 3.9868 ≈ 4.0

From the first and second entries, the order of the reaction with respect to ClO- is 1. The third and fourth entries give fractional orders. Check the data.

From practical experience, it seems that the order of the reaction with respect to ClO- is 1.0. Put the value of the rate and [ClO-] and a = 1 in the rate law to get k.

k = Rate/[ClO-]a = (0.0015618 M/s)/(0.152 M) = 0.010275 s-1

The rate law for the reaction is

Rate = (0.010275 s-1)*[ClO-]